# Simulation-Based Inference for Probabilistic Galaxy Detection and Deblending

### v1

@@ 3.3 Detection and classification encoders
BLISS uses forward amortized variational inference (FAVI) (ambrogioni2019forward) to perform posterior inference for light source properties other than galaxy parameters. Two of the BLISS encoders are optimized using FAVI: the detection and classification encoders. As presented in Equations 2 through 4, the detection encoder outputs the parameter of a Bernoulli distribution describing whether a source is present in a given tile. It also outputs two parameters describing where the source centroid is within the tile. The classification encoder outputs a parameter indicating whether a source is a star or a galaxy.
We now describe our procedure for training these encoders. In this section, we refer to the corresponding light source parameters captured by the detection and classification encoders as $\mathcal{Z}^{\prime}$, where $\mathcal{Z}^{\prime}$ is specific to a given image with $T$ tiles:
$\mathcal{Z}^{\prime}\equiv\{n_{t},b_{t},\bm{\ell}_{t}\}_{t=1}^{T}.$ (7)
@@
$L(\bm{\phi})=\mathbb{E}_{\mathcal{Z}\sim P(\mathcal{Z}|\mathcal{X})}D_{\rm KL}(P(\mathcal{Z}|\mathcal{X})\,\|\,Q_{\bm{\phi}}(\mathcal{Z})).$ (8)
Here, $D_{\rm KL}(\cdot\,\|\,\cdot)$ denotes the KL divergence, which, for generic distributions $p$ and $q$, is
$D_{\rm KL}(p\,\|\,q)\equiv\int p(z)\log\left(\frac{q(z)}{p(z)}\right)dz.$ (9)
@@
$L^{\prime}(\bm{\phi})=\mathbb{E}_{\mathcal{X,Z}\sim P(\mathcal{X,Z})}\sum_{t=1}^{T}$ $\Big[\log Q_{\bm{\phi}}(n_{t})+\log Q_{\bm{\phi}}(\bm{\ell}_{t}|n_{t})$
@@
Finally, we optimize both neural networks with a stochastic gradient procedure using the Adam optimizer (kingma2014adam). We use a learning rate of $10^{-4}$ and otherwise default parameters of the PyTorch implementation (pytorch2019). We use mini-batches of size $32$ from dataset-blend. [^13] We chose this combination of parameters because doing so leads to a smooth and rapid decline in the loss function curve during training. The hyperparameters of other encoders in this work are chosen similarly. The neural network architecture for these encoders follows a standard ResNet-like architecture (e.g., he2016deep) consisting of convolutions, batch norms, dropouts, and ReLUs. [^14] The exact implementation of our neural network can also be found in our public repository: https://github.com/ismael-mendoza/bliss.
@@ 3.4 Deblending encoder
We train the last of the BLISS encoders, the deblending encoder, using a two-step approach that is similar to the procedure in arcelin2021deblending.
@@
Finally, we emphasize the distinct statistical approaches to detection/classification and inferring galaxy representations. In the detection and classification encoders, each light-source parameter is associated with a variational distribution factor that captures the uncertainty in that parameter. In contrast, the deblending encoder returns only point estimates with no uncertainty. Despite this, we can construct posterior catalogs on galaxy properties that capture detection uncertainty by using samples from the detection encoder as inputs to the deblending encoder. This procedure will be demonstrated in Section 4.6. Future work could use a VAE (as in arcelin2021deblending) or FAVI (patel2025npe) to capture the full uncertainty in the properties of blended galaxies.
@@ 4.1 SNR and Blendedness
We use the aperture photometry module in SEP (barbary2016sep) to measure the signal-to-noise ratio (SNR) defined as
@@
We adopt the definition of blendedness from bosch2018hyper. For a galaxy with intensity $\bm{I}$, blendedness is
@@ 4.2 Detection evaluation
We evaluate the trained detection encoder on galaxy blends and compare it with SEP, a Python implementation of SourceExtractor (bertin1996sextractor; barbary2016sep). Specifically, we look at the precision, recall, and $F_{1}$ score of the detection encoder using different probability thresholds for what constitutes a detection. For the BLISS detection encoder, for each tile having a detection probability larger than a given threshold, we choose the centroid prediction to be the mean of the predicted centroid posterior on that tile. For SEP, we use the configuration described in Appendix C, which is designed to match a previous blending study (sanchez2021effects). We do not attempt to vary the SEP parameters.
We first define matched detections as detections that are within $2$ pixels of a true centroid and assigned to a unique true centroid using the Hungarian matching algorithm (kuhn1955hungarian) (available as the Scipy function linear_sum_assignment). This is a bipartite graph matching algorithm that minimizes a certain cost function on the graph edges. The cost function is
@@ 4.3 Classification evaluation
We additionally use the precision, recall, and $F_{1}$-Score metrics, as in the previous section, to evaluate the encoder. These metrics are computed differently for this classification context. First, we split these metrics into a galaxy version and a star version, as shown in Figure 5. This is helpful as dataset-blend is highly imbalanced, with a much larger number of galaxies than stars. We say that the model classifies a light source as a galaxy if the output probability of the source being a galaxy is >0.5>0.5, and a star otherwise. [^17] The optimal choice for this probability cut will depend on the specific application. The performance demonstrated in Figure 5 can be tuned by modifying this value as needed. The precision for galaxies (stars) in this context corresponds to the number of correctly classified galaxies (stars) divided by the total number of sources the model classified as galaxies. The recall for galaxies (stars) is then the number of correctly classified galaxies (stars) divided by the total number of galaxies. The $F_{1}$ score is computed as in Equation 17. These quantities are binned and plotted in Figure 5 based on the true SNR of the sources. We exclude the faintest sources with $\text{SNR}<10$ from this figure. In the left plot, we can see that all of the galaxy metrics are above $90\%$ for all relevant SNRs. This is expected given our highly unbalanced dataset. The precision decreases rapidly from around $98\%$ to $93\%$ at the SNR roughly corresponding to the vertical strip in Figure 4. The recall increases as the SNR decreases as the model defaults to the prior and classifies most objects as galaxies. Meanwhile, we see that all the star metrics are monotonic as a function of SNR. This is consistent with Figure 4 as the purple dots consistently move from the top to the bottom as the SNR increases.
Our model becomes more uncertain when classifying low SNR sources, as expected. However, for very low SNR, the prior dominates, and a significant number of stars are classified as galaxies with high confidence. One way to make progress could be to incorporate additional prior information into the network that takes into account the model of stars (i.e., the PSF). For example, using the spread model estimator (e.g., mohr2012spread).
@@ 4.6 Joint Detection and Deblending Uncertainty
Our results are summarized as a function of the blendedness of the central galaxy in Figure 9. We choose a total of $20$ blendedness bins, where $18$ of them are in the range $3.80\times 10^{-4},3.76\times 10^{-1}]$ and are set to ensure an equal number of objects in each. Objects with values lower than $3.76\times 10^{-4}$ or higher than $3.56\times 10^{-1}$ are placed in the first and last bin, respectively. The upper bound was chosen to ensure enough statistics on high blendedness examples. Each bin contains at least $400$ corresponding images used to obtain residuals. The curves represent the (fractional) median residual in each blendedness bin from each method, where the “Samples” approach uses the average of the flux measured across the samples to compute its residual. The shaded regions correspond to the $1\sigma$ bootstrap errors on the median obtained by resampling the residuals on each bin. In this figure, we see that the sampling approach to measure photometry significantly outperforms the point estimations from both “MAP” and “SEP” approaches across blendedness bins. The gap in the median residuals increases with blendedness and is particularly large for the last bin, where the residual jumps from approximately $16\%$ to about $160\%$ for the MAP and about $400\%$ for SEP.
@@
We hypothesize that these improvements are driven by the large number of objects undetected by the “MAP” or “SEP” approach when the blendedness level is high or the SNR is low. Undetected objects will not be removed from the image through deblending, and would thus lead to flux overestimation. On the other hand, the sampling approach has the potential to detect more sources as no hard cutoff is used for detection probability. Samples corresponding to sources in the image that are difficult to detect could, on average, reduce the flux residual. It is also possible that the sampling approach is less sensitive to the tile boundary artifact, as the probability of a source close to a boundary appears to be shared between the corresponding tiles (liu2023variational). However, the comparison using SEP detections indicates that this cannot be the sole source of improvement, as the SEP algorithm does not employ a tiling procedure.
@@ 5 Discussion and Summary
We use an autoencoder (AE) trained on parametric simulated galaxies as our galaxy model. This model enables the training of the deblending encoder and the reconstruction of galaxies (Section 3.4). This approach follows a growing body of work on using deep neural networks for galaxy image modeling. Recent work on using deep learning for galaxy modeling commonly uses generative adversarial networks (GANs) and variational autoencoders (VAEs), and, more recently, score-based models (ravanbakhsh2016enabling; dia2019galaxy; fussell2019forging; reiman2019deblending; lanusse2020deep; arcelin2021deblending; hemmati2022deblending; smith2022realistic; biswas2024madness; sampson2024score). We used an AE in BLISS since autoencoders are generally more stable in training than GANs (kingma2019introduction) and have been shown to perform galaxy image reconstruction accurately (lanusse2020deep; arcelin2021deblending). lanusse2020deep used VAEs to build deep generative models that can recreate realistic galaxy images using data from the HST COSMOS catalog (mandelbaum2018HSC). VAEs have also been used successfully for deblending (arcelin2021deblending). These types of methods have the potential to mitigate the model bias often incurred by more standard parametric galaxy models, as demonstrated by remy2022model. Currently, we do not fully leverage the modeling capabilities of deep generative models in BLISS, as our AE is deterministic and is trained exclusively on parametric galaxies. However, we plan to incorporate realistic galaxy profiles and use a VAE to capture the full uncertainty of galaxy properties in future work.
@@
BLISS includes a detection encoder (Section 3.3) that is capable of outputting a posterior distribution for the centroids of stars and galaxies in astronomical images. Traditional detection approaches are based on peak finding or thresholding, such as the popular algorithm SExtractor (bertin1996sextractor; barbary2016sep). However, classical approaches to cataloging cannot characterize uncertainty in the number of sources found and are known to result in a significant number of ambiguously blended sources (dawson2016ambiguous). Probabilistic approaches for detection relying on MCMC have been proposed for deblending crowded starfields, such as in feder2020multiband, and predicting point estimates and uncertainties of star fluxes and centroids. However, these approaches are computationally expensive because they rely on MCMC. In this work, we evaluated BLISS point (deterministic) predictions of source detections and found comparable performance with the chosen SourceExtractor configuration (Section 4.2). However, in this study, we discarded the probabilistic information from the BLISS detection encoder by using thresholds. In a later section, we show an example of how this additional information can be used (Section 4.6).
@@
BLISS also includes a classification encoder (Section 3.3) that outputs a probability that a detected source is either a galaxy or a star. Approaches to star-galaxy separation have thus far often involved neural networks, such as the one used in SExtractor (bertin1996sextractor). This neural network differs from ours in that it does not use the images directly but rather 10 carefully chosen numerical features and a feedforward architecture. More recent approaches to star-galaxy classification include convolutional neural networks such as BLISS and avoid the need to extract features carefully (kim2016star; garg2022star). Other approaches have used low-dimensional embeddings of optical images and Gaussian processes for classification (goumiri2020star; muyskens2022star). The CNN architecture used in BLISS could enable the use of information from multiple bands or exposures for a given light source. These approaches will become more important as new surveys observe fainter objects, where the increase in blending of smaller galaxies makes classification difficult (slater2020morphological).
@@
BLISS performs deblending using its deterministic deblending encoder (Section 3.4), which takes as input samples from the detection and classification encoders. The deblending encoder consists of a neural network that learns to target the galaxy at the center of an image, while removing the flux of surrounding galaxies. This approach to deblending is similar to DebVAdEr (arcelin2021deblending) and MADNESS (biswas2024madness). Other deblenders, such as Scarlet (melchior2018scarlet), use constrained optimization to deblend light sources without deep learning. In future work, we also plan to conduct an extensive comparison of BLISS with these and other deblenders using the BlendingToolKit package (mendoza2025blending).
@@
As explored in Section 4.5, our tiling procedure (Section 3.1) has potential drawbacks. We found that detection probability correlates strongly with distance from the source to the tile boundary. This makes the detection probability drop below $0.5$ even for detectable and isolated sources. This effect could be impacting the results of Section 4.2, degrading detections and reducing the BLISS recall across all SNR and blendedness bins. In a precursor paper to BLISS, which uses a similar detection encoder architecture (liu2023variational), the authors find that the probability is “shared” between tiles when the source becomes too close to the tile boundary. This suggests that drawing samples from the detection encoder might be more robust to the tile boundary effect, as all tiles with significant assigned probabilities can contribute. Other ways to mitigate this artifact include allowing neighboring tiles to overlap, or modifying the variational distribution (Equation 1) to encode dependencies between neighboring tiles (regier2025autoregressive). Future work will involve a more in-depth analysis of tile boundary effects and exploring these and other mitigation strategies.
@@
In Section 4.6, we quantified the effect on photometric measurements of combining the probabilistic detection and deblending capabilities of BLISS. We generate aperture flux posterior distributions of individual deblended galaxies by repeatedly sampling galaxy centroids from the detection encoder and using these samples as input to the deterministic deblender encoder. These flux posterior distributions primarily capture the contribution from the detection uncertainty of the target and nearby sources. This procedure could be applied to produce posteriors for any measurable galaxy property, although we only explored improvements to aperture fluxes in this work. Future work will replace the deblender encoder with a VAE to better approximate the full uncertainty present in blended galaxy systems. We find that using these deblended flux posteriors can drastically reduce the median flux residual for highly blended galaxies in our slightly simplified dataset-central simulations (Figure 9). While out of scope for this paper due to our use of single-band image simulations, future work using multi-band images could investigate downstream applications to improve color and redshift estimation by deriving their corresponding posteriors. Previous work has already shown that a machine learning deblender can be used to improve photometric redshift estimation in simulations (merz2025deepdisc).
@@
Future work on BLISS will focus on preparing this algorithm for use with real astronomical data and supporting cosmological analyses from stage-IV optical surveys such as LSST. Specifically, we plan to: (1) Extend BLISS so that processing multi-band data becomes possible. This might enable propagating our predicted flux posteriors (Section 4.6) to photo-z estimation. (2) Accommodate data with spatially variable PSFs (patel2025npe) and background, as well as other effects present in real survey images. (3) Relatedly, training and testing our models using more realistic LSST-like image simulations such as the LSST DESC DC2 simulations (abolfathi2021dc2; duan2025bliss_dc2) and the OpenUniverse 2024 simulations (openuniverse2024). (4) Training and testing our encoders using images of galaxies with realistic morphology. For instance, we could consider combining deep fields and space observations from other cosmological surveys to build a more realistic model of galaxy morphology. In this work, we have used only parametric galaxies, which may introduce model biases in real survey data. (5) Test and validate our method using a specific set of science-driven metrics. For example, use BLISS detection samples to explore whether shear detection biases could be mitigated. (6) Address potential biases related to the tiling procedure and corresponding independence assumptions in BLISS (Section 4.5). This includes the fact that we allow at most one source centroid per tile. (7) Understand how robust BLISS is to significant and relevant differences between the training and test data. (8) Finally, develop a general methodology that can be used to propagate BLISS predicted uncertainties to downstream cosmological analyses.

### v2

@@ 3.3 Detection and classification encoders
BLISS uses forward amortized variational inference (FAVI) (Ambrogioni et al., 2019) to perform posterior inference for light source properties other than galaxy parameters. Two of the BLISS encoders are optimized using FAVI: the detection and classification encoders. As presented in Equations 2 through 4, the detection encoder outputs the parameter of a Bernoulli distribution describing whether a source is present in a given tile. It also outputs two parameters describing where the source centroid is within the tile. The classification encoder outputs a parameter indicating whether a source is a star or a galaxy.
We now describe our procedure for training these encoders. In this section, $\mathcal{Z}$ refers to the light source parameters captured by the detection and classification encoders, so that:
$\mathcal{Z}\equiv\{n_{t},b_{t},\bm{\ell}_{t}\}_{t=1}^{T}.$ (7)
@@
where $\mathbb{E}$ denotes the expectation operator [^11] The expectation operator is defined by $\mathbb{E}_{\bm{x}\sim P(\bm{x})}f=\int f(\bm{x})P(\bm{x})\,d\bm{x}$, for a generic function $f$ and a random variable $\bm{x}$ with probability distribution $P$. and $D_{\rm KL}(\cdot\,\|\,\cdot)$ the KL divergence, which, for generic distributions $p$ and $q$, is
$D_{\rm KL}(p\,\|\,q)\equiv\int p(z)\log\left(\frac{p(z)}{q(z)}\right)dz.$ (9)
@@
$L^{\prime}(\bm{\phi})=-\mathbb{E}_{\mathcal{X,Z}\sim P(\mathcal{X,Z})}\sum_{t=1}^{T}$ $\Big[\log Q_{\bm{\phi}}(n_{t})+\log Q_{\bm{\phi}}(\bm{\ell}_{t}|n_{t})$
@@
Finally, we optimize both neural networks with a stochastic gradient procedure using the Adam optimizer (Kingma and Ba, 2014). We use a learning rate of $10^{-4}$ and otherwise default parameters of the PyTorch implementation (Paszke et al., 2019). We use mini-batches of size $32$ from dataset-blend. [^14] We chose this combination of parameters because doing so leads to a smooth and rapid decline in the loss function curve during training. The hyperparameters of other encoders in this work are chosen similarly. The neural network architecture for these encoders follows a standard ResNet-like architecture (e.g., He et al., 2016) consisting of convolutions, batch norms, dropouts, and ReLUs. [^15] The exact implementation of our neural network can also be found in our public repository: https://github.com/prob-ml/bliss/tree/desc-oja.
@@ 3.4 Deblending encoder
We train the last of the BLISS encoders, the deblending encoder, using a two-step approach that is similar to the procedure in Arcelin et al. (2021).
@@
Finally, we emphasize the distinct statistical approaches to detection/classification and inferring galaxy representations. In the detection and classification encoders, each light-source parameter is associated with a variational distribution factor that captures the uncertainty in that parameter. In contrast, the deblending encoder returns only point estimates with no uncertainty. Despite this, we can construct posterior catalogs on galaxy properties that capture detection uncertainty by using samples from the detection encoder as inputs to the deblending encoder. This procedure will be demonstrated in Section 4.6. Future work could use a VAE (as in Arcelin et al., 2021) or FAVI (Patel et al., 2025) to capture the full uncertainty in the properties of blended galaxies.
@@ 4.1 SNR and Blendedness
We use the aperture photometry module in SEP, a Python implementation of SourceExtractor (Bertin and Arnouts, 1996; Barbary, 2016), to measure the signal-to-noise ratio (SNR) defined as
@@
We adopt the definition of blendedness from Bosch et al. (2018). For a galaxy with intensity $\bm{I}$, blendedness is
@@ 4.2 Detection evaluation
We evaluate the trained detection encoder on galaxy blends and compare it with SEP. Specifically, we look at the precision, recall, and $F_{1}$ score of the detection encoder using different probability thresholds for what constitutes a detection. For the BLISS detection encoder, for each tile having a detection probability larger than a given threshold, we choose the centroid prediction to be the mean of the predicted centroid posterior on that tile. For SEP, we use the configuration described in Appendix C, which is designed to match a previous blending study (Sanchez et al., 2021a). We do not attempt to vary the SEP parameters.
We first define matched detections as detections that are within $2$ pixels of a true centroid and assigned to a unique true centroid using the Hungarian matching algorithm (Kuhn, 1955) (available as the Scipy function linear_sum_assignment). This is a bipartite graph matching algorithm that minimizes a certain cost function on the graph edges. The cost function is
@@ 4.3 Classification evaluation
We additionally use the precision, recall, and $F_{1}$-Score metrics, as in the previous section, to evaluate the encoder. These metrics are computed differently for this classification context. First, we split these metrics into a galaxy version and a star version, as shown in Figure 5. This is helpful as dataset-blend is highly imbalanced, with a much larger number of galaxies than stars. We say that the model classifies a light source as a galaxy if the output probability of the source being a galaxy is >0.5>0.5, and a star otherwise. [^18] The optimal choice for this probability cut will depend on the specific application. The performance demonstrated in Figure 5 can be tuned by modifying this value as needed. The precision for galaxies (stars) in this context corresponds to the number of correctly classified galaxies (stars) divided by the total number of sources the model classified as galaxies. The recall for galaxies (stars) is then the number of correctly classified galaxies (stars) divided by the total number of galaxies. The $F_{1}$ score is computed as in Equation 17. These quantities are binned and plotted in Figure 5 based on the true SNR of the sources, and bins are constructed so that they contain the same number of galaxies or stars. We exclude the faintest sources with $\text{SNR}<10$ from this figure for visualization purposes. In the left plot, we can see that all of the galaxy metrics are above $90\%$ for all relevant SNRs. This is expected given our highly unbalanced dataset. The precision decreases rapidly from around $98\%$ to $93\%$ at the SNR roughly corresponding to the vertical strip in Figure 4. The recall increases as the SNR decreases as the model defaults to the prior and classifies most objects as galaxies. Meanwhile, we see that all the star metrics are monotonic as a function of SNR. This is consistent with Figure 4 as the purple dots consistently move from the top to the bottom as the SNR increases.
Our model becomes more uncertain when classifying low SNR sources, as expected. However, for very low SNR, the prior dominates, and a significant number of stars are classified as galaxies with high confidence. One way to make progress could be to incorporate additional prior information into the network that takes into account the model of stars (i.e., the PSF). For example, using the spread model estimator (e.g., Mohr et al., 2012).
@@ 4.6 Joint Detection and Deblending Uncertainty
Our results are summarized as a function of the blendedness of the central galaxy in Figure 9. We choose a total of $20$ blendedness bins, where $18$ of them are in the range $3.80\times 10^{-4},3.76\times 10^{-1}]$ and are set to ensure an equal number of objects in each. Objects with values lower than $3.80\times 10^{-4}$ or higher than $3.76\times 10^{-1}$ are placed in the first and last bin, respectively. The upper bound was chosen to ensure enough statistics on high blendedness examples. Each bin contains at least $400$ corresponding images used to obtain residuals. The curves represent the (fractional) median residual in each blendedness bin from each method, where the “Samples” approach uses the average of the flux measured across the samples to compute its residual. The shaded regions correspond to the $1\sigma$ bootstrap errors on the median obtained by resampling the residuals on each bin. In this figure, we see that the sampling approach to measure photometry significantly outperforms the point estimations from both “MAP” and “SEP” approaches across blendedness bins. The gap in the median residuals increases with blendedness and is particularly large for the last bin, where the residual jumps from approximately $16\%$ to about $160\%$ for the MAP and about $400\%$ for SEP.
@@
We hypothesize that these improvements are driven by the large number of objects undetected by the “MAP” or “SEP” approach when the blendedness level is high or the SNR is low. Undetected objects will not be removed from the image through deblending, and would thus lead to flux overestimation. On the other hand, the sampling approach has the potential to detect more sources as no hard cutoff is used for detection probability. Samples corresponding to sources in the image that are difficult to detect could, on average, reduce the flux residual. It is also possible that the sampling approach is less sensitive to the tile boundary artifact, as the probability of a source close to a boundary appears to be shared between the corresponding tiles (Liu et al., 2023). However, the comparison using SEP detections indicates that this cannot be the sole source of improvement, as the SEP algorithm does not employ a tiling procedure.
@@ 5 Discussion and Summary
We use an autoencoder (AE) trained on parametric simulated galaxies as our galaxy model. This model enables the training of the deblending encoder and the reconstruction of galaxies (Section 3.4). This approach follows a growing body of work on using deep neural networks for galaxy image modeling. Recent work on using deep learning for galaxy modeling commonly uses generative adversarial networks (GANs) and variational autoencoders (VAEs), and, more recently, score-based models (Ravanbakhsh et al., 2016; Dia et al., 2019; Fussell and Moews, 2019; Reiman and Göhre, 2019; Lanusse et al., 2020; Arcelin et al., 2021; Hemmati et al., 2022; Smith et al., 2022; Biswas et al., 2024; Sampson et al., 2024). We used an AE in BLISS since autoencoders are generally more stable in training than GANs (Kingma and Welling, 2019) and have been shown to perform galaxy image reconstruction accurately (Lanusse et al., 2020; Arcelin et al., 2021). Lanusse et al. (2020) used VAEs to build deep generative models that can recreate realistic galaxy images using data from the HST COSMOS catalog (Mandelbaum et al., 2018). VAEs have also been used successfully for deblending (Arcelin et al., 2021). These types of methods have the potential to mitigate the model bias often incurred by more standard parametric galaxy models, as demonstrated by Remy et al. (2022). Currently, we do not fully leverage the modeling capabilities of deep generative models in BLISS, as our AE is deterministic and is trained exclusively on parametric galaxies. However, we plan to incorporate realistic galaxy profiles and use a VAE to capture the full uncertainty of galaxy properties in future work.
@@
BLISS includes a detection encoder (Section 3.3) that is capable of outputting a posterior distribution for the centroids of stars and galaxies in astronomical images. Traditional detection approaches are based on peak finding or thresholding, such as the popular algorithm SExtractor (Bertin and Arnouts, 1996; Barbary, 2016). However, classical approaches to cataloging cannot characterize uncertainty in the number of sources found and are known to result in a significant number of ambiguously blended sources (Dawson et al., 2015). Probabilistic approaches for detection relying on MCMC have been proposed for deblending crowded starfields, such as in Feder et al. (2020), and predicting point estimates and uncertainties of star fluxes and centroids. However, these approaches are computationally expensive because they rely on MCMC. In this work, we evaluated BLISS point (deterministic) predictions of source detections and found comparable performance with the chosen SourceExtractor configuration (Section 4.2). However, in this study, we discarded the probabilistic information from the BLISS detection encoder by using thresholds. In a later section, we show an example of how this additional information can be used (Section 4.6).
@@
BLISS also includes a classification encoder (Section 3.3) that outputs a probability that a detected source is either a galaxy or a star. Approaches to star-galaxy separation have thus far often involved neural networks, such as the one used in SExtractor (Bertin and Arnouts, 1996). This neural network differs from ours in that it does not use the images directly but rather 10 carefully chosen numerical features and a feedforward architecture. More recent approaches to star-galaxy classification include convolutional neural networks such as BLISS and avoid the need to extract features carefully (Kim and Brunner, 2016; Garg et al., 2022). Other approaches have used low-dimensional embeddings of optical images and Gaussian processes for classification (Goumiri et al., 2020; Muyskens et al., 2022). The CNN architecture used in BLISS could enable the use of information from multiple bands or exposures for a given light source. These approaches will become more important as new surveys observe fainter objects, where the increase in blending of smaller galaxies makes classification difficult (Slater et al., 2020).
@@
BLISS performs deblending using its deterministic deblending encoder (Section 3.4), which takes as input samples from the detection and classification encoders. The deblending encoder consists of a neural network that learns to target the galaxy at the center of an image, while removing the flux of surrounding galaxies. This approach to deblending is similar to DebVAdEr (Arcelin et al., 2021) and MADNESS (Biswas et al., 2024). Other deblenders, such as Scarlet (Melchior et al., 2018), use constrained optimization to deblend light sources without deep learning. In future work, we also plan to conduct an extensive comparison of BLISS with these and other deblenders using the BlendingToolKit package (Mendoza et al., 2025).
@@
As explored in Section 4.5, our tiling procedure (Section 3.1) has potential drawbacks. We found that detection probability correlates strongly with distance from the source to the tile boundary. This makes the detection probability drop below $0.5$ even for detectable and isolated sources. This effect could be impacting the results of Section 4.2, degrading detections and reducing the BLISS recall across all SNR and blendedness bins. In a precursor paper to BLISS, which uses a similar detection encoder architecture (Liu et al., 2023), the authors find that the probability is “shared” between tiles when the source becomes too close to the tile boundary. This suggests that drawing samples from the detection encoder might be more robust to the tile boundary effect, as all tiles with significant assigned probabilities can contribute. Other ways to mitigate this artifact include allowing neighboring tiles to overlap, or modifying the variational distribution (Equation 1) to encode dependencies between neighboring tiles (Regier, 2025). Future work will involve a more in-depth analysis of tile boundary effects and exploring these and other mitigation strategies.
@@
In Section 4.6, we quantified the effect on photometric measurements of combining the probabilistic detection and deblending capabilities of BLISS. We generate aperture flux posterior distributions of individual deblended galaxies by repeatedly sampling galaxy centroids from the detection encoder and using these samples as input to the deterministic deblender encoder. These flux posterior distributions primarily capture the contribution from the detection uncertainty of the target and nearby sources. This procedure could be applied to produce posteriors for any measurable galaxy property, although we only explored improvements to aperture fluxes in this work. Future work will replace the deblender encoder with a VAE to better approximate the full uncertainty present in blended galaxy systems. We find that using these deblended flux posteriors can drastically reduce the median flux residual for highly blended galaxies in our slightly simplified dataset-central simulations (Figure 9). While out of scope for this paper due to our use of single-band image simulations, future work using multi-band images could investigate downstream applications to improve color and redshift estimation by deriving their corresponding posteriors. Previous work has already shown that a machine learning deblender can be used to improve photometric redshift estimation in simulations (Merz et al., 2025).
@@
Future work on BLISS will focus on preparing this algorithm for use with real astronomical data and supporting cosmological analyses from stage-IV optical surveys such as LSST. Specifically, we plan to: (1) Extend BLISS so that processing multi-band data becomes possible. This might enable propagating our predicted flux posteriors (Section 4.6) to photo-z estimation. (2) Accommodate data with spatially variable PSFs (Patel et al., 2025) and background, as well as other effects present in real survey images. (3) Relatedly, training and testing our models using more realistic LSST-like image simulations such as the LSST DESC DC2 simulations (LSST-DESC et al., 2021; Duan et al., 2025) and the OpenUniverse 2024 simulations (OpenUniverse et al., 2025). (4) Training and testing our encoders using images of galaxies with realistic morphology. For instance, we could consider combining deep fields and space observations from other cosmological surveys to build a more realistic model of galaxy morphology. In this work, we have used only parametric galaxies, which may introduce model biases in real survey data. (5) Test and validate our method using a specific set of science-driven metrics. For example, use BLISS detection samples to explore whether shear detection biases could be mitigated. (6) Address potential biases related to the tiling procedure and corresponding independence assumptions in BLISS (Section 4.5). This includes the fact that we allow at most one source centroid per tile. (7) Understand how robust BLISS is to significant and relevant differences between the training and test data. (8) Finally, develop a general methodology that can be used to propagate BLISS predicted uncertainties to downstream cosmological analyses.